OpenStudy (anonymous):

solve: x+y=13 4x+2y=13 by substitution or elimination technique

OpenStudy (slaaibak):

Get one variable in terms of the other one: x=13 - y Now plug this x into 4x + 2y = 13 it becomes: 4(13 - y) + 2y = 13 Now just solve it.

OpenStudy (anonymous):

I solved to get the x I cant find y

OpenStudy (slaaibak):

Once you find x, plug it into x + y = 13 to get the y value

OpenStudy (anonymous):

thanks got x=9 y=4

OpenStudy (slaaibak):

I think you went wrong somewhere. I'll help a bit: By expanding 4(13 - y) + 2y = 13 you get: 52 - 4y + 2y = 13 -2y = -39 y=19.5 x= 13 - 19.5 x= -6.5

OpenStudy (anonymous):

this is where I got the equations so I cant get negative numbers. In a farmyard there are only chickens and sheep. There are 13 animals altogether and between them they have 34 legs. Use simultaneous equations to determine how many chickens and sheep there are.

OpenStudy (slaaibak):

x + y = 13 4x + 2y = 34 You just made a mistake by letting the second equation equal 13 and not 34

OpenStudy (anonymous):

oh yeap sorry